AnimeGhoul8863:

4. Create a unique parabola in the pattern f(x) = ax2 + bx + c. Describe the direction of the parabola and determine the y-intercept and zeros.

AnimeGhoul8863:

My equation o G(x)=x3−3x2−4x+12

AnimeGhoul8863:

@Ultrilliam

AnimeGhoul8863:

@BenLindquist

BenLindquist:

@Shadow

AnimeGhoul8863:

@Shadow

563blackghost:

Did you make the equation? Parabola format is \(\bf{ax^{2}+bx+c}\) you started with \(\bf{x^{3}}\) this makes the equation into a cube format. How about setting it at just \(\bf{-3x^{2}-4x+12}\)?

AnimeGhoul8863:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AnimeGhoul8863 My equation o G(x)=x3−3x2−4x+12 \(\color{#0cbb34}{\text{End of Quote}}\)

AnimeGhoul8863:

so if we took ^^^ and did it as f(x)=-3x^2-4x+12

AnimeGhoul8863:

that would be the parabola

563blackghost:

yes

563blackghost:

now you can find the zeroes.

563blackghost:

To find it it's best to use the quadratic formula. \(\Large\bf{x=\frac{-b \pm \sqrt{b^{2}-4ac}}{2a}}\) Just plug in the equation. `a = -3` `b = -4` `c=12` So you would have to solve for 2. \(\Large\bf{x=\frac{-(-4) + \sqrt{(-4)^{2}-4(-3)(12)}}{2(-3)}}\) \(\Large\bf{x=\frac{-(-4) - \sqrt{(-4)^{2}-4(-3)(12)}}{2(-3)}}\) This will give you your zeroes.

AnimeGhoul8863:

im sooooooo confused what are the zeros

AnimeGhoul8863:

@563blackghost

563blackghost:

If you don't know on how to find the zeroes, then I can not help you since the equation is asking for info you have not learned. @AnimeGhoul8863 Here is a site that can help find the zeroes. http://www.biology.arizona.edu/biomath/tutorials/quadratic/roots.html

AnimeGhoul8863:

@563blackghost im not saying i cant find the zeros im saying im a little confused on ur equations but ok

563blackghost:

What are you exactly confused on? There are 2 ways you can find the roots, squaring the function or using the quadratic formula. I simply used the quadratic formula.

AnimeGhoul8863:

i guess im confused bc im not good at fractions or formula's